### v1

@@ IV-C2 Ablation Study on Reward Function
Figure 8: Results of ablation study on reward function. The 5 tasks are opening/closing drawer, opening/closing laptop and turning faucet. The task is considered to be failed if it is not done when time step reaches 50.
IV-D Experiments on Dexterous Hand
IV-D1 Real World Articulated Object Manipulation
For each 3 categories, we choose one object for real object manipulation experiments. Considering the FOV of the RGBD camera as well as the workspace of motion planning, we randomly set the location and initial joint status $s_{0}$ of articulated objects on the table in a certain range, such that the object is in the workspace of the manipulator. We randomly select $\Delta s_{target}$ which does not exceed the joint limit and covers both directions of possible movement.
For each category, we conduct 30 experiments. For parameters of Sampling-based Model Predictive Control module, we find that $T=50$, $N=100$, $h=10$ leads to fast searching as well as good performance. For the parameters in the reward function, we make adjustments based on the results of simulation experiments. We set $\omega_{s}=20$, $\omega_{t}=50$, $\omega_{contact}=10$, $\omega_{d}=10$, $\omega_{c}=0.001$ and $\omega_{v}=0.01$. We use eigen dimension $m=2$ to conduct real world manipulation. We use 10 processes for sampling in simulation on a computer that has an Intel Core i7-12700 CPU and an NVIDIA 3080Ti GPU. It takes about 2.5 minutes to find a feasible trajectory.
Results of all the experiments can be found in Fig. 9, and statistical results can be found in Table II. Trajectories of opening and closing a drawer are shown in Fig. 10. Similar to manipulation with 2-finger gripper, the drawer has the lowest $\left|\delta_{r}\right|$ and the faucet has the highest $\left|\delta_{r}\right|$ according to Table II.
@@ IV-D1 Real World Articulated Object Manipulation
TABLE II: Accuracy of real articulated objects manipulation with dexterous hand
Category Drawer Laptop Faucet
Number of manipulations 32 31 26
<20% 14 12 4
<40% 22 23 18
Number of manipulations s.t. $\left|\delta_{r}\right|$ <60% 27 27 22
Avg $\left|\delta\right|$ 1.90cm 6.92 9.48
Avg $\left|\delta_{r}\right|$ 28.25% 30.76% 45.72%
@@
IV-D2 Ablation Studies
For the ablation study of dexterous hand manipulation, we investigate the impact of several key factors. Specifically, we analyze the influences of eigengrasp dimensions on Sampling-based Model Predictive Control, study how pixel projection affects the Interactive Perception module and also explore the influences of different reward functions.
Eigengrasp Dimension. In Section III-C2, we propose to utilize eigengrasp to reduce the dimension of action space, enhancing the efficiency of our search process. To evaluate the effectiveness of MPC with different eigengrasp dimensions, we conduct experiments with dimensions $m=1,2,7,16$. We adopt 3 performance metrics: success rate, joint jerk, and computation time. We evaluate on 5 different tasks in simulation: opening/closing laptop, turning faucet, and opening/closing drawer. Each task is repeated 10 times with randomized object positions and initial robot configurations.
@@ IV-D2 Ablation Studies
    Success rate Fig.11 presents the success rate of each task with different dimensions. The tasks of opening laptop and closing drawer achieve a 100% success across all the 4 dimension numbers. This high success rate likely stems from the simplicity of these tasks. However, in the tasks of closing the laptop and opening the drawer, trajectories generated with $m=1$ shows a significantly lower success rate compared to the other dimensions. This may be attributed to the reduced dexterity of the Allegro Hand when operating in a 1-DOF configuration. Interestingly, for all the 5 tasks, the dimension reduced to $m=2$ performs comparably to $m=7,16$. This finding highlights the effectiveness of using eigengrasp space with $m=2$, achieving similar success rates while reducing computation time.
@@
    Joint jerk For real robot control, joint jerk reflects the smoothness of a robot’s movements. Our study investigates how the eigengrasp dimension affects the jerks of dexterous hand joints. To quantify jerk, we utilize the third derivative of hand joint positions with respect to time. Specifically, we calculate the average jerk for each finger at every time step, reflecting the smoothness of finger movement. Fig 12 illustrates the results for the 5 tasks. Our experimental results consistently demonstrate that finger joint jerks increase with higher eigengrasp dimensions. Remarkably, for all the 5 tasks, dimension $m=1$ exhibits superior performance in terms of thumb jerk compared to $m=2,7,16$. However, for the remaining three fingers: both $m=1,2$ demonstrate similar advantages over higher dimensions.
@@
Pixel Projection. In Section III-A2, we propose a pixel projection method that leverages both RGB images and depth information of an object to transforms a 2D post-contact vector into a 3D robot trajectory. To evaluate the necessity of the pixel projection approach, we compare it with randomly generated vectors based solely on 2D affordance. Specifically, we select one object per category and generate three random vectors for each object. The results, shown in Fig. 14, demonstrate that the vector synthesized by pixel transformation is better suited for executing one-step interactions compared to the randomly generated direction vector.
Reward Function. In Section III-C2, we design reward function in the sampling-based model predictive control module for dexterous hand manipulation tasks. Five terms of reward functions are designed, which include $r_{success}$, $r_{target}$, $r_{contact}$, $r_{dist}$ and $r_{reg}$. We conduct 5 groups of experiments to reveal each term’s influence against the full reward function. The first group runs iCEM with the full reward function as in III-C. The other 4 experiments drop one term of the full reward function. To make the ablation result more generalizable, we conduct 3 tasks for each group: opening laptop, closing laptop and turning faucet. In each task, we randomize the position of the object, the initial joint angle of the robot as well as the target qpos of the object to generate 30 different trajectory running iCEM. The task that is not done when time step reaches 50 is considered failed. Fig. 15 and Fig. 15 respectively summarize the experimental results.
@@
IV-D3 Advantage of Dexterous Manipulation
@@
Figure 17: Trajectory comparison of a dexterous hand and a 2-finger gripper performing the same laptop-closing task. The dexterous hand completes the task in fewer steps.
IV-E Effectiveness of Interactive Perception
The Interactive Perception module is designed to improve the accuracy of the constructed world model by utilizing the two different point clouds captured before and after interaction. To evaluate its necessity, we train another model using a single-frame point cloud as the network input. For each category of objects, we select one real object and compare the modeling results of two-frame and single-frame point cloud inputs. Fig. 18 shows the comparison results. The findings demonstrate that actively interacting with the movable part of the object and altering its state allows us to build a transition model with more accurate segmentation of movable parts and joint axis estimation, which is necessary for precise manipulation.
Figure 18: Ablation study on Interactive Perception. The model constructed with Interactive Perception has more accurate movable part segmentation and joint axis estimation.
IV-F Advanced Manipulation Skills
@@
V Conclusion and Discussion
@@
For future work, we plan to integrate proprioceptive sensing and tactile sensing during real-robot interaction to refine the constructed world model for more precise manipulation. 3D generative AI has seen great progress in the last few years. We also plan to integrate the AIGC technique to improve the geometry quality of the digital twin. Besides, a module that estimates the state of the object in real time will enhance reactive manipulations. Lastly, we aim to expand the framework to include mobile manipulation, objects with multiple movable parts and deformable objects, thereby broadening its applicability across various robotic tasks.

### v2

@@ IV-D2 Ablation Study on Reward Function
Figure 11: Results of ablation study on reward function. The 4 tasks are opening/closing drawer and opening/closing laptop. The task is considered to be failed if it is not done when time step reaches 50.
IV-E Experiments on Dexterous Hand
IV-E1 Real World Articulated Object Manipulation
We choose two categories: Drawer and Cabinet, and conduct 20 experiments for each category. Considering the FOV of the RGBD camera as well as the workspace of motion planning, we randomly set the location and initial joint status $s_{0}$ of articulated objects on the table in a certain range, such that the object is in the workspace of the manipulator. We randomly select $\Delta s_{target}$ which does not exceed the joint limit and covers both directions of possible movement.
We use the Allegro Hand to conduct all the quantitative experiments of dexterous hand. For parameters of Sampling-based Model Predictive Control module, we find that $T=50$, $E=100$, $h=10$ leads to fast searching as well as good performance. For the parameters in the reward function, we make adjustments based on the results of simulation experiments. We set $\omega_{s}=20$, $\omega_{t}=50$, $\omega_{contact}=10$, $\omega_{d}=10$, $\omega_{c}=0.001$ and $\omega_{v}=0.01$. We use eigen dimension $m=2$ to conduct real world manipulation. We use 30 processes for sampling in simulation on a computer that has an AMD Ryzen 9 9950X 16-Core CPU and an NVIDIA 3080Ti GPU. It takes about 3 minutes to find a feasible trajectory.
Results of all the experiments can be found in Fig. 12, and statistical results can be found in Table III. Trajectories of opening and closing a drawer are shown in Fig. 13.
@@ IV-E1 Real World Articulated Object Manipulation
TABLE III: Accuracy of real articulated objects manipulation with dexterous hand
Category Drawer Cabinet
Number of manipulations 20 20
<10% 3 3
<30% 11 18
Number of manipulations s.t. $\left|\delta_{r}\right|$ <50% 20 20
Avg $\left|\delta\right|$ 1.64cm 6.87
Avg $\left|\delta_{r}\right|$ 26.4% 20.4%
@@
IV-E2 Ablation Studies
For the ablation study of dexterous hand manipulation, we investigate the influences of eigengrasp dimensions on Sampling-based Model Predictive Control the influences of different reward functions.
Eigengrasp Dimension
In Section III-C, we propose to utilize eigengrasp to reduce the dimension of action space, enhancing the efficiency of our search process. To evaluate the effectiveness of MPC with different eigengrasp dimensions, we conduct experiments with dimensions $m=1,2,7,16$. We adopt 3 performance metrics: success rate, joint jerk, and computation time. We evaluate on 4 different tasks in simulation: opening/closing laptop and opening/closing drawer. Each task is repeated 10 times with randomized object positions and initial robot configurations.
@@ IV-E2 Ablation Studies
    Success rate Fig.14 presents the success rate of each task with different dimensions. The tasks of opening laptop and closing drawer achieve a 100% success across all the 4 dimension numbers. This high success rate likely stems from the simplicity of these tasks. However, in the tasks of closing the laptop and opening the drawer, trajectories generated with $m=1$ shows a significantly lower success rate compared to the other dimensions. This may be attributed to the reduced dexterity of the Allegro Hand when operating in a 1-DOF configuration. Interestingly, for all the 4 tasks, the dimension reduced to $m=2$ performs comparably to $m=7,16$. This finding highlights the effectiveness of using eigengrasp space with $m=2$, achieving similar success rates while reducing computation time.
@@
    Joint jerk For real robot control, joint jerk reflects the smoothness of a robot’s movements. Our study investigates how the eigengrasp dimension affects the jerks of dexterous hand joints. To quantify jerk, we utilize the third derivative of hand joint positions with respect to time. Specifically, we calculate the average jerk for each finger at every time step, reflecting the smoothness of finger movement. Fig 15 illustrates the results for the 4 tasks. Our experimental results consistently demonstrate that finger joint jerks increase with higher eigengrasp dimensions. Remarkably, for all the 4 tasks, dimension $m=1$ exhibits superior performance in terms of thumb jerk compared to $m=2,7,16$. However, for the remaining three fingers: both $m=1,2$ demonstrate similar advantages over higher dimensions.
@@
Reward Function
In Section III-C3, we design reward function in the sampling-based model predictive control module for dexterous hand manipulation tasks. Five terms of reward functions are designed, which include $r_{success}$, $r_{target}$, $r_{contact}$, $r_{dist}$ and $r_{reg}$. We conduct 5 groups of experiments to reveal each term’s influence against the full reward function. The first group runs iCEM with the full reward function as in III-C3. The other 4 experiments drop one term of the full reward function. To make the ablation result more generalizable, we conduct 2 tasks for each group: opening laptop and closing laptop. In each task, we randomize the position of the object, the initial joint angle of the robot as well as the target qpos of the object to generate 30 different trajectory running iCEM. The tasks that are not done when time step reaches 50 are considered failed. Fig. 17 and Fig. 17 respectively summarize the experimental results.
@@
IV-E3 Advantage of Dexterous Manipulation
@@
Figure 19: Trajectory comparison of a dexterous hand and a 2-finger gripper performing the same drawer opening task. The dexterous hand completes the task in fewer steps (17 vs. 22).
IV-F Experiments on Explicit World Model Construction
IV-F1 Model Accuracy Evaluation
In this section, we quantitatively evaluate the model construction accuracy and compare it with ScrewNet [19]. We use 4 metrics: RDE (Relative Depth Error), which is the relative error between the rendered depth map of the constructed model and the real depth map captured by the RGBD camera; IoU (Intersection over Union), which measures the degree of alignment between the rendered mask and the groundtruth mask; Angle Err and Pos Err, which measure the joint axis error. For prismatic joints, only the angle error is applicable. Since the training labels and the predictions of ScrewNet are defined in the object frame, which is determined by the CAD model’s frame definition, we only evaluate the angle error of the joint axis for ScrewNet.
We conduct experiments on all the four categories and collect 10 pairs of frames for each category. Table IV presents the evaluation results. Our method demonstrates the ability to accurately construct the world models across all four categories, achieving an average angle error of less than $9^{\circ}$. The relatively low IoU for lamps can be attributed to the small area the lamp occupies in the image, making it more sensitive to reconstruction errors.
TABLE IV: Quantitative evaluation of explicit world model construction accuracy
Category RDE IoU Pos Err Ang Err Ang Err of
(%) (m) () ScrewNet [19] ()
Laptop 1.47 0.92 0.022 1.34 23.15
Cabinet 3.77 0.94 0.029 4.84 11.67
Drawer 2.23 0.94 - 8.58 23.21
Lamp 2.12 0.80 0.009 7.95 9.38
IV-F2 Impact of Model Construction Accuracy on Manipulation Performance
In this section, we evaluate the impact of model construction accuracy on manipulation performance. We introduce different scales of joint position errors and joint orientation errors into the constructed model, search for the trajectories, and execute these trajectories on the real robot. Fig. 20 shows the manipulation results. All the experiments are conducted on Laptop. As can be seen, the manipulation error is positively correlated with the model construction error, with position errors exerting a more significant impact.
Figure 20: Real-world manipulation results of model construction with different scales of joint position errors and joint orientation errors.
IV-F3 Ablation Study on Movable Part Segmentation
In this section, we validate the impact of movable part segmentation on kinematic structure construction. In [47], segmentation is jointly optimized with part motions using a distance loss. However, it may produce incorrect segmentation results when only two or three frames of point clouds are used. To address this, our proposed pipeline first segments the meshes and then estimates the kinematic structure. Table V compares the results across all four categories. Our method achieves more accurate segmentation and improved axis parameter estimation compared to [47], particularly for prismatic joints. Fig. 21 provides a visualization of some results. [47] may generate an incorrect number of parts and wrong joint types.
TABLE V: Comparison of model construction accuracy with and without our movable part segmentation.
Category Success Angle Err () $\downarrow$ Pos Err (m) $\downarrow$
Laptop w/ Part 10/10 1.34 0.022
w/o Part 9/10 1.32 0.048
Cabinet w/ Part 10/10 4.84 0.029
w/o Part 10/10 6.10 0.034
Drawer w/ Part 10/10 8.58 -
w/o Part 5/10 16.04 -
Lamp w/ Part 10/10 7.95 0.009
w/o Part 9/10 7.54 0.015
Figure 21: Visualization of model construction results: (a)(c)with our movable part segmentation, (b)(d)without our movable part segmentation. Revolute joints are shown in pink, and prismatic joints are shown in blue.
IV-G Ablation Studies on Affordance Detection
IV-G1 Where2Act VS. VRB
Fig. 22 shows the affordance detection results of Where2Act and VRB across four categories. Since the goal of the one-step interaction is to change the state of the movable part, a pushing interaction is sufficient in most cases. For objects where pushing is not feasible—such as fully closed cabinets and drawers, the affordance detection module must detect the handles. For cabinets, VRB can detect handles in 17 out of 20 samples, while Where2Act can detect 7. For drawers, VRB can detect handles in 3 out of 20 samples, and Where2Act can only detect 1. VRB achieves a much higher recall rate than Where2Act and generates more correct interaction directions for the following two reasons: First VRB is trained on human manipulation videos sourced from the real-world domain, directly aligning with the robot’s operational environment, whereas Where2Act is trainiend using simulation data. Second, the scale of available human manipulation videos significantly exceeds that of 3D interactable assets, enabling better generalizability. The reason the recall rate for drawer handles is inferior is that the handles of the drawers used in our work are very small. As a result, they occupy only a small number of pixels in both images and point clouds, making their detection particularly challenging for the two methods. In real-world experiments, we first visualize the predicted action of Where2Act and VRB in the 3D space and manually choose the better one to execute. In over 90% of scenarios, VRB generates better action than Where2Act.
Figure 22: Visualization of affordance detection results: results of VRB are shown in green arrows, and results of Where2Act are shown in red arrows.
IV-G2 Pixel Projection
In Section III-A2, we propose a pixel projection method that leverages both RGB images and depth information of an object to transforms a 2D post-contact vector into a 3D robot trajectory. To evaluate the necessity of the pixel projection approach, we compare it with randomly generated vectors based solely on 2D affordance. Specifically, we select one object per category and generate three random vectors for each object. The results, shown in Fig. 23, demonstrate that the vector synthesized by pixel transformation is better suited for executing one-step interactions compared to the randomly generated direction vector.
IV-H Objects With Multiple Movable Parts
In this section, we demonstrate the effectiveness of our method on objects with multiple movable parts. During the interactive perception stage, our method sequentially changes the states of each movable part. Then in the explicit world construction stage, we first separate the generated meshes into movable parts, construct the kinematic structure, and estimate the axis direction and origin of each joint. Fig. 24 illustrates the modeling results of our method applied to two different objects. As shown, our approach accurately reconstructs the explicit models of objects with varying kinematic structures.
Figure 24: Multi-frame observations before and after interactions of objects with multiple movable parts and reconstructed articulated models with axes and texture: (a)(b) A cabinet with two revolute joints; (c)(d) A fridge with two revolute joints (shown in pink) and one prismatic joint(shown in blue).
IV-I Advanced Manipulation Skills
@@
V Analysis and Discussion
In this section, we analyze the reason of manipulation errors and discuss the limitations of our method.
By comparing the manipulation outcomes in simulation and real-world objects, we conclude that the manipulation errors can be attributed to the following factors, in addition to model construction errors:
  1. (1)
    The dynamic properties of the real articulated objects are complicated. For example, the elastic deformation of laptops, that the laptop partially returns to its original shape once the contact force from the robot is removed, is not modeled in the simulation.
  2. (2)
    The kinematic structure of a real articulated object is not ideal. For example, there might be gaps in the drawer rails, which transform an ideal prismatic joint into a joint with multiple DoFs.
Our explicit world model construction module infers kinematic structures by analyzing inter-frame differences, which necessitates observable state changes in object parts during the interactive perception stage. Another limitation is that for novel or uncommon objects, existing 3D AIGC techniques may produce meshes that are not perfectly aligned with real-world observations, affecting the accuracy of the resulting digital twins. However, with the rapid advancements in AIGC techniques, we anticipate these alignment challenges will be effectively addressed in the near future.
VI Conclusion